### v3

# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

Abstract
The rapid advancement of large language models (LLMs) has revolutionized role-playing, enabling the development of general role-playing models. However, current role-playing training has two significant issues: (I) Using a predefined role profile to prompt dialogue training for specific scenarios usually leads to inconsistencies and even conflicts between the dialogue and the profile, resulting in training biases. (II) The model learns to imitate the role based solely on the profile, neglecting profile-dialogue alignment at the sentence level. In this work, we propose a simple yet effective framework called Beyond Dialogue, designed to overcome these hurdles. This framework innovatively introduces “beyond dialogue” tasks to align dialogue with profile traits based on each specific scenario, thereby eliminating biases during training. Furthermore, by adopting an innovative prompting mechanism that generates reasoning outcomes for training, the framework allows the model to achieve fine-grained alignment between profile and dialogue at the sentence level. The aforementioned methods are fully automated and low-cost. Additionally, the integration of automated dialogue and objective evaluation methods forms a comprehensive framework, paving the way for general role-playing. Experimental results demonstrate that our model excels in adhering to and reflecting various dimensions of role profiles, outperforming most proprietary general and specialized role-playing baselines. All code and datasets are available at https://github.com/yuyouyu32/BeyondDialogue.
Introduction
The rapid advancement of large language models (LLMs) has demonstrated their significant potential in user interactions (Achiam et al. 2023; Bai, Bai, and et al. 2023; Dubey et al. 2024). A particularly promising area is the development of role-playing LLM agents, which are capable of simulating both real and fictional roles to deliver immersive and interactive experiences (Chen et al. 2024b, a). Applications like Character AI (Shazeer and Freitas 2024), Replika (Replika 2024), and Baichuan-NPC (Baichuan 2024) exemplify the growing impact of these agents. However, the proprietary nature of these platforms restricts users’ ability to customize specific role models. Thus, the development of an open-source, general role-playing model is imperative. Such a model provides a reliable foundation for expressing predefined role profiles and customizing role models across scenarios (Tencent 2024; Baichuan 2024).
Figure 1: An example of the bias issue between predefined role profile and dialogues in role-playing training.
At the current stage of role-playing dialogue training tasks, role profiles are typically manually compiled (Shao et al. 2023) or generated through LLM summaries (Zhou et al. 2023), including information such as personality and speaking style as depicted throughout a whole novel or script (Zhou et al. 2023; Li et al. 2023). However, a bias exists between the training dialogues extracted from individual scenarios and the predefined role profiles. As an illustration in Figure 1, Hermione’s speaking style is predefined into four categories, which are manually extracted from the entire content of the novel. However, the training dialogues for the scenario in Figure 1 predominantly reflect Hermione’s bookish (blue) and encouraging (green) styles. This bias arises because the predefined profile is generally not fully represented in a single scenario, leading to inconsistencies where the prompted role does not align with, or even contradicts, the training dialogues. Such biases are common: in the HPD dataset (Chen et al. 2023), we sampled 800 scenarios and found that 69.68% of the dialogues showed a bias from the predefined character profiles, while 50.69% exhibited a bias in style representation. These inconsistencies between predefined role profiles and the training dialogue corpus, as notably evident in our §Main Results experiments, can mislead the model, hindering its ability to accurately reflect the role profile during dialogues. Meanwhile, the model has only learned a vague mapping from profile to dialogue, without fine-grained alignment of specific traits to their corresponding dialogues. As shown in Figure 1, the model fails to map the traits “bookish” and “encouraging” to the specific dialogue sentences highlighted in blue and green. Consequently, the model is unable to learn how “personality traits are revealed in dialogues”.
These observations highlight two critical issues in the current role-playing training process:
  • Bias Between Profile and Dialogues: A dialogue corpus that deviates from preset profiles may introduce bias during training, impairing the model’s ability to follow the profiles.
  • Inadequate Fine-Grained Alignment: A single dialogue training task fails to fine-grained align role dialogues with their profiles, as it lacks the nuanced understanding of how specific traits manifest, limiting the model’s ability to comprehend and represent the roles’ complex traits.
We propose a new training role-playing framework—Beyond Dialogue to overcome the aforementioned issues and advance toward general role-playing. We incorporate powerful LLM through a prompting mechanism approach to align role profiles with scenario role-playing dialogues. This not only ensures the reliability of the LLM’s reasoning but also explicitly creates a fine-grained alignment dataset at the sentence level. Taking inspiration from actors learning to play different roles—understanding the performance of various role traits in scenarios to enhance their portrayal (Zarrilli 2005)—we also employ fine-grained alignment tasks to train the role-playing model. These beyond-dialogue alignment tasks help the model understand how to represent specific role traits at the sentence level, thereby enhancing its overall effectiveness in role portrayal.
During the evaluation process, we assess the model’s role-playing capability by examining how well the dialogue aligns with the defined role profiles, which is consistent with the beyond-dialogue alignment tasks. Traditional evaluation methods have largely depended on subjective assessments, such as human evaluations or LLMs providing consistent scores between dialogue and profiles. These approaches are often irreproducible and lead to significant evaluation variance. Consequently, we have converted all evaluation tasks into objective tasks, with the evaluation objective clearly defined as the model’s ability to follow the role profiles. To ensure both efficiency and reproducibility in the evaluation, we employ an approach that combines automatic dialogues with the “LLMs as Judger” (Kim et al. 2023) method to assess different role-playing models.
We applied Beyond Dialogue to bilingual chat LLM baselines—Qwen2 (Bai, Bai, and et al. 2023) and Mistral (Jiang et al. 2023)—to empirically evaluate its effectiveness. Extensive experiments demonstrate that Beyond Dialogue significantly enhances the ability of LLMs to follow and portray role profiles. Under our framework, the trained models surpassed existing proprietary advanced general baselines, such as GPT-4o (OpenAI 2024), and specialized proprietary role-playing baselines, like Baichuan-NPC-Turbo (Baichuan 2024), in key dimensions of general role-playing performance. Moreover, the beyond dialogue alignment tasks consistently improved the models’ role-playing evaluation accuracy by over 30%, approaching the evaluation capabilities of GPT-4o. This alignment supports the development of a self-generating, self-evaluating, and self-aligning pipeline for role-playing models, fostering continuous improvement in their performance.
Our main contributions are as follows:
  1. 1.
    We identified a bias in role-playing training where the predefined role’s profile does not align with the dialogues at the specific scenario scope. Furthermore, we demonstrated that this bias significantly hinders the model’s ability to effectively follow the predefined profile.
  2. 2.
    We propose Beyond Dialogue, a novel general role-playing training framework that employs LLMs with a prompting mechanism to align role profiles with scenario-specific dialogues. This approach derives fine-grained alignment tasks at the sentence level, which are incorporated into the role-playing training, further enhancing the effectiveness of role portrayal.
  3. 3.
    We introduce a novel evaluation pipeline that converts all assessment tasks into objective tasks, focusing on the model’s ability to follow user-defined role profiles. By combining automatic dialogues with the “LLMs as Judger” method, our approach ensures both efficiency and reproducibility in role-playing model evaluation.
Related Work
Role-Playing Dataset/Framework Dataset Automatically Constructed? Scenario Style Character Personality Emotion Relationship Evaluation Scope
LIGHT (Urbanek et al. 2019) Sentence
PDP (Han et al. 2022) Sentence
ChatHaruhi (Li et al. 2023) Sentence
HPD (Chen et al. 2023) Sentence
CharacterLLM (Shao et al. 2023) Multi-Turn Artificial Interview
CharacterGLM (Zhou et al. 2023) Multi-Turn Artificial Dialogue
RoleLLM (Wang et al. 2023c) Sentence
CharacterEval (Tu et al. 2024) Sentence
RoleCraft (Tao et al. 2023) Sentence
InCharacter (Wang et al. 2023a) Multi-Turn Artificial Interview
DITTO (Lu et al. 2024) Multi-Turn Automated Dialogue
Beyond Dialogue Multi-Turn Automated Scenario Dialogue
Table 1: Comparison of Beyond Dialogue with other datasets or framework. “✓” indicates that the dataset/framework meets the given standard or includes the specified dimension in the role profile, while “✗” indicates that it does not. “Sentence” indicates that the dataset/framework evaluates performance based on the final sentence of multi-turn dialogues. Other methods assess performance through dialogues/interviews generated manually or automatically.
General Role-playing Agents. Recent research on general role-playing agents has largely overlooked the issue of bias between predefined profiles and scenario-specific dialogues. Instead, the focus has been on constructing role-playing datasets, primarily through manual curation (Zhou et al. 2023), extraction from novels (Li et al. 2023; Tao et al. 2023), or generation using GPT (Shao et al. 2023; Wang et al. 2023c), to acquire extensive annotated role-playing data. We compared various current Role-Playing Datasets/Frameworks and summarized five key dimensions for constructing role profiles and evaluations in Table 1: Style, Character, Personality, Emotion, and Relationship. Additionally, scenarios provide the background and context for dialogues, making the role’s conversations more contextual and realistic (Chen et al. 2023; Shao et al. 2023).
Role-playing Evaluation. The initial step in evaluating role-playing models involves generating dialogues for assessment, employing three primary methods: Providing Historical Dialogues (Tu et al. 2024; Wang et al. 2023c; Chen et al. 2023; Tao et al. 2023), Manual Dialogue (Zhou et al. 2023; Wang et al. 2023a; Shao et al. 2023), and Automated Dialogue with LLMs (Lu et al. 2024). The first method is limited to evaluating the model within the scope of individual sentences; however, single sentences are likely to fail to fully capture the entire role profile and may be biased by exposure to novels or scripts used during pre-training. In contrast, the second and third methods can generate complete scenario dialogues, which are more conducive to comprehensive evaluations.
There are three common methods for evaluating dialogues generated by models: Metric-based Evaluation (Li et al. 2023; Wang et al. 2023c), Human Evaluation (Zhou et al. 2023), and “LLMs as Judges” (Shao et al. 2023; Wang et al. 2023c; Lu et al. 2024). The first method primarily assesses the model’s retention of the standard response, which is derived from the original text of the novel or manually annotated compilation. The second method, while more accurate, is costly and difficult to replicate, limiting its broader application. The third method, however, is attracting increasing attention due to its high efficiency, low cost, and scalability.
Hybrid Task Training LLMs. In recent years, hybrid task training has emerged as a pivotal method for fine-tuning LLMs. For instance, the LLaMA Writing Assistant is trained on a variety of textual tasks (Wang et al. 2023b). Incorporating tasks like summarization, text generation, and reasoning within reading comprehension has greatly improved the model’s language understanding and QA abilities (Cheng, Huang, and Wei 2023). Furthermore, training with a combination of general and role-specific instructions has improved the model’s role-playing abilities (Tao et al. 2023). These studies demonstrate that incorporating tasks directly or indirectly related to downstream tasks in hybrid training can significantly enhance model performance.
Beyond Dialogue Framework
Figure 2: Illustration of Beyond Dialogue Training Framework and Automated Dialogue Evaluation. The left side shows the training phases, which include role-playing dialogue, chit-chat, and profile alignment. The profile alignment results are utilized to adjust each scenario’s dialogue profiles, eliminating training biases. On the right, the LLM generates random scenarios and roles for multi-turn dialogues with the model, followed by an evaluation using objective questions to obtain quantitative metrics of the model’s role-playing capabilities.
As shown in Figure 2, In the Beyond Dialogue training framework, role profiles are aligned and adjusted for each scenario to create ’pure’ role-playing dialogue data, which is then combined with alignment reasoning and chit-chat data for model training. In the Automated Dialogue Evaluation, the pipeline generates random scenarios and multi-turn dialogues to interact with the model, using objective questions to quantify its role-playing capabilities.
The design and implementation details of each component will be elaborated in the subsequent chapters.
Alignment Dataset Construction
Figure 3: An schematic of our automated pipeline for constructing Beyond Dialogue dataset.
To advance general role-playing, we require a diverse set of role settings and dialogue data. Comparing novel-extracted dialogues (Chen et al. 2023) with those generated by GPT models (Wang et al. 2023c; Li et al. 2023) shows that the latter lacks human likeness, while the former better captures role nuances. However, datasets that meet the requirement for diverse roles, extracted solely from novels, are not open-source due to high costs and significant labor. Therefore, we propose a low-cost (as discussed in §E), fully automated pipeline for constructing a role-playing dialogue dataset (Figure 3) and introduce a method to address biases and fine-grained alignment issues mentioned in the §Introduction:
  1. 1.
    Role-playing Dialogue Dataset Construction: We first segment the text and apply a role frequency threshold to filter chunks. Given the large initial chunk volume, we employ open-source models to extract scenarios and evaluate chunks, retaining only those that reflect role traits and belong to single scenes. Next, we compare various LLMs and ultimately select GPT-4o for dialogue extraction and validation, ensuring the dialogues are coherent and free from conflicts with role profiles. By applying rules and open-source models to filter all chunks, we generate high-quality chunks for proprietary model processing. This approach enables the low-cost, fully automated construction of a high-quality role-playing dialogue dataset, laying a solid foundation for subsequent derivative tasks and model training.
  2. 2.
    Profile Alignment: After constructing the dataset, we use GPT-4o to align each multi-turn dialogue across five dimensions: Character, Style, Emotion, Relationship, and Personality (CSERP). Using an innovative prompting mechanism (as detailed in §F), GPT-4o provides detailed reasoning processes, such as identifying how a dialogue line reflects a specific speaking style. This approach generates fine-grained CSERP alignment tasks, helping the model maintain its reasoning capabilities and enhancing its ability to perceive and adhere to profiles.
  3. 3.
    Profile Adjustment: Based on the alignment results, we dynamically adjust the Profile settings for each dialogue to ensure consistency. This process addresses the common issue of bias between scene dialogues and role profiles in role-playing datasets, aligning the training inputs (Profile prompts) with outputs (dialogue labels), thereby ensuring effective training.
All prompts used for dataset construction and alignment can be found in §F. A detailed explanation of the database construction process is provided in §A, along with the statistical information about our dataset, obtained using this method, which can be found in §A. This section also includes manually collected and organized role profile statistics, covering the five dimensions of CSERP. A discussion of the low cost is provided in §E.
Supervised Finetuning
Using the constructed dataset, we fine-tuned LLMs with a mix of Role-Playing Dialogue data $D_{r}$, Deriving CSERP data $D_{d}$, and Chit-Chat data $D_{c}$ in a 1:5:4 ratio. $D_{r}$ dataset was aligned and adjusted to ensure consistency between role profiles and dialogue across scenarios. The learning objective was to minimize the aggregate negative log-likelihood across all datasets, as described by:
$\min_{\pi}\mathbb{E}_{(s,a)\sim D_{r}\odot D_{d}\odot D_{c}}\left[-\log\pi(a|s)\right]$
where $s$ is the prompt input, and $a$ is the corresponding label, $\odot$ represents hybrid task training, and $\pi$ denotes the trained LLMs. The total volume of training data, consistent with other comparative experiments, was regulated to be ten times that of the benchmark role-playing data, effectively managing the potential impact of data volume on LLM training. Specific training parameters are detailed in §D.
Automated Dialogue Evaluation
Evaluating the effectiveness and reproducibility of role-playing is still an open problem. As discussed in §Related Work, the diversity in dialogue generation and evaluation methods complicates establishing widely accepted standards.
Our another contribution is to propose an automated dialogue method, as shown on the right side of Figure 2. We first generate a role and its description, aligning it with the worldview of the role to be evaluated. Based on their profiles, we create a dialogue scenario, design the emotions, and define the relationship between the roles. Finally, the two models engage in multi-turn dialogues within this context, producing a dialogue corpus for subsequent role-playing model evaluations.
During the evaluation phase, we adopted the “LLMs as judges” (Kim et al. 2023) approach, designing five tasks based on the role profile dimensions as discussed in §Related Work: Character, Style, Emotion, Relationship, and Personality. Besides, we used Human-likeness to assess whether the model’s output matches human expression and Coherence to evaluate dialogue continuity. Additionally, we introduced a role-based multiple-choice evaluation to assess the model’s role consistency across multi-turn dialogues(Lu et al. 2024). We reformatted all evaluation tasks into multiple-choice and true/false questions, clarifying criteria, streamlining the process, and reducing biases. This approach has a high level of acceptance in manual checks and closely mirrors human evaluation, making the evaluation more systematic and reliable for large-scale applications. All prompts used for dialogue generation and evaluation are available in §F.
Experiments
Experimental Setup
Dataset. Following the method outlined in §Alignment Dataset Construction, we created the RP dataset by extracting 280 Chinese and 31 English roles from 123 novels and scripts, resulting in 3,552 scenario dialogues with 23,247 turns, all based on authentic sources. Compared to the other datasets listed in §A Table 7, ours has the highest number of real roles and the highest turns with real dialogues, which is significant for building a general role-playing model. Additionally, we created the CSERP dataset by deriving five alignment training tasks from each of the 3,552 sessions, resulting in five training data points per session. After aligning and adjusting role profiles with their dialogues, the data was transformed into the RPA dataset. The CC dataset included both Chinese (NaturalConv (Wang et al. 2021)) and English (DailyDialog (Li et al. 2017)) chit-chat data.
Baselines. We test both open-source and proprietary advanced bilingual (Chinese-English) chatbots in our assessment framework. Our proprietary general baselines include: GPT-4o, GPT-3.5-Turbo (OpenAI 2024), Yi-Large-Turbo (Wanwu 2024), Deepseek-Chat (Deepseek 2024). These models represent the current state-of-the-art in language generation technology, covering different architectures and training methods. Additionally, we selected several baselines focused on role-playing: Index-1.9B-Character (bilibili 2024) (open-source), Baichuan-NPC-Turbo (Baichuan 2024) (proprietary), and CharacterGLM (Zhou et al. 2023) (open-source). These models are specifically designed and optimized for role-playing, capable of generating more realistic character dialogues and interactions. Furthermore, to validate the effectiveness of our framework, we trained two open-source bilingual baselines: Qwen2-7B-Instruct (Bai, Bai, and et al. 2023) and Mistral-Nemo-Instruct-2407 (Jiang et al. 2023).
Metrics Design. As previously mentioned, our goal is to evaluate the model’s ability to follow the role’s profile and scenario setting, which aligns with the objectives of the derivative tasks of CSERP discussed in §A Figure 5. Consequently, the evaluation tasks and alignment tasks under the five dimensions—Character, Style, Emotion, Relationship, and Personality—are identical. Furthermore, we introduced three additional evaluation metrics: Human-likeness, Coherence, and Role Choice of the dialogues:
  • Emotion and Relationship: Emotion is assessed based on Ekman’s six basic emotions (Ekman 1992), and both Emotion and Relationship are rated by GPT-4o on a scale of 0-10 based on the dialogue being evaluated. The scoring results are used to calculate the Normalized Mean Absolute Percentage Error (NMAPE) with the context labels generated by the prompt model.
  • Character, Style, and Personality: Rated by GPT-4o based on the dialogue to be evaluated, referencing the candidate set of labels provided by the role profile. Notably, Personality uses the MBTI classification system and is a binary recall task, whereas Character and Style are multi-label recall tasks.
  • Human-likeness: Used to evaluate whether the role-playing model can provide a more natural and realistic interaction experience. We convert the evaluation of human-likeness into a true/false question. With few-shot prompting, GPT-4o determines whether the input dialogue is human-generated or model-generated.
  • Role Choice: Used to evaluate the recognizability of the role presentation in the dialogue. We mask the names in the dialogue and have GPT-4o select the most suitable role for the current dialogue from the four candidate roles provided.
  • Coherence: Used to evaluate the logical consistency and contextual coherence of the multi-turn dialogues. GPT-4o assesses whether the dialogue is coherent based on the given context and scenario.
Main Results
Model Character Style Emotion Relationship Personality Avg. ↑ Human -likeness ↑ Role Choice ↑ Coherence ↑
Recall ↑ Recall ↑ NMAPE ↓ NMAPE ↓ Precision ↑
General Baselines(Proprietary)
GPT-4o 74.32 ± 1.15 81.67 ± 1.51 16.31 ± 0.48 12.13 ± 0.66 66.58 ± 4.41 78.83 ± 1.64 67.33 ± 3.95 87.33 ± 3.86 99.67 ± 0.33
GPT-3.5-Turbo 72.26 ± 1.27 73.66 ± 1.73 17.79 ± 0.56 14.17 ± 0.73 66.92 ± 4.85 76.18 ± 1.83 $3$ 4.43 83.00 ± 4.68 97.33 ± 1.17
Yi-Large-Turbo 75.13 ± 1.22 79.18 ± 1.58 16.44 ± 0.49 13.48 ± 0.67 68.25 ± 4.61 78.53 ± 1.72 47.00 ± 4.60 84.33 ± 3.67 92.67 ± 2.39
Deepseek-Chat 75.46 ± 1.14 81.49 ± 1.51 15.92 ± 0.46 12.42 ± 0.63 67.92 ± 4.57 79.30 ± 1.66 52.33 ± 4.95 83.00 ± 4.68 96.67 ± 1.00
Role-play Expertise Baselines
Index-1.9B-Character 73.33 ± 1.32 76.48 ± 1.50 17.99 ± 0.53 13.58 ± 0.71 66.33 ± 4.57 76.92 ± 1.73 21.67 ± 3.96 78.67 ± 5.14 69.67 ± 3.85
CharacterGLM-6B 73.36 ± 1.28 76.08 ± 1.55 18.58 ± 0.55 14.27 ± 0.79 67.33 ± 4.34 76.79 ± 1.70 16.00 ± 2.38 81.00 ± 4.40 25.67 ± 3.48
Baichuan-NPC-Turbo 75.19 ± 1.23 79.15 ± 1.38 17.24 ± 0.51 13.10 ± 0.69 65.33 ± 4.84 77.87 ± 1.73 56.00 ± 4.66 86.33 ± 4.90 99.00 ± 0.56
Custom Trained Baselines
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 77.53 ± 1.61 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
+ RP & CC 71.56 ± 1.26 74.66 ± 1.51 17.36 ± 0.49 12.58 ± 0.69 68.17 ± 4.32 76.89 ± 1.65 53.67 ± 3.76 86.33 ± 4.22 86.33 ± 2.42
+ RPA & CC 74.44 ± 1.14 77.63 ± 1.40 16.74 ± 0.46 12.07 ± 0.67 69.50 ± 4.31 78.55 ± 1.60 56.00 ± 3.41 85.00 ± 4.49 91.67 ± 1.80
+ RPA & CC & CSERP 74.58 ± 1.28 78.47 ± 1.45 16.62 ± 0.48 11.38 ± 0.67* 69.08 ± 4.46 78.83 ± 1.67 59.00 ± 4.46 87.00 ± 4.73 92.67 ± 1.59
[…]
+ RP & CC 74.91 ± 1.21 78.59 ± 1.39 16.52 ± 0.48 12.28 ± 0.67 67.00 ± 4.31 78.34 ± 1.61 56.67 ± 3.85 82.00 ± 4.90 90.00 ± 1.59
+ RPA & CC 76.43 ± 1.18 81.28 ± 1.37 16.10 ± 0.45 12.35 ± 0.67 67.83 ± 4.36 79.42 ± 1.61 62.33 ± 3.21 87.33 ± 3.77 92.67 ± 1.85
+ RPA & CC & CSERP 78.67 ± 1.12* 82.52 ± 1.33* 15.68 ± 0.5* 11.22 ± 0.72* 69.67 ± 4.27 80.79 ± 1.59* 64.33 ± 3.80* 87.33 ± 3.74 99.00 ± 0.56
Table 2: Main results of Beyond Dialogue. We report the average scores with their standard error of the mean (SEM). Bold numbers indicate the highest scores, while underlined numbers are the best in the group, * means the improvements over the untrained baseline are statistically significant (achieving $p<0.05$ in t-test). The Avg. column represents the mean of the Character, Style, Emotion, Relationship, and Personality metrics, which are directly defined within the profile, with Emotion and Relationship calculated as 1-NMAPE. All dialogues are evaluated by GPT-4o as judges.
We conducted a total of 300 bilingual evaluations for each model, with each evaluation consisting of a scenario containing five turns of dialogue. Specific evaluating parameters are detailed in §D. The evaluation process was fully automated, with GPT-4o generating chat roles and scenarios, and engaging in multiple turns of dialogue with the evaluated model. The generated dialogues were input into an automatic evaluation pipeline to obtain quantitative performance metrics for the role-playing models. These evaluation results were used to calculate the average performance of each model across five dimensions and three additional metrics.
We present the main results in Table 2. Due to the lack of role-playing training, the role-playing capabilities of the general baselines are closely related to their general capabilities. In the Human-likeness dimension, dialogues generated by GPT-4o have a 67.33% probability of being perceived as real human dialogues, significantly outperforming other general models. This is crucial for creating authentic dialogue scenario simulations.
Among the role-playing expertise baselines, the open-source models Index-1.9B-Character and characterGLM-6B, which were fine-tuned using only role-playing data, showed significantly lower performance in Human-likeness and Coherence compared to other models. In contrast, the proprietary model Baichuan-NPC-Turbo demonstrated a significant improvement in Human-likeness.
In our experiments, the model trained with unaligned role-playing data (+RP & CC) exhibited training biases, leading to limited or even diminished performance in the unaligned dimensions of Character, Style, and Personality. However, the model trained with aligned data (+RPA & CC) showed significant improvements in the character and style dimensions, with average scores surpassing both general-purpose and expertise baselines, indicating the importance of aligned data in mitigating training bias.
Under the Beyond Dialogue framework, the fully enhanced Qwen-2-7B models obtained through supervised fine-tuning achieved the highest scores in most dimensions, with its human-likeness significantly outperforming other baselines except GPT-4o, and showing statistically significant improvements across five dimensions compared to the untrained baseline ($p<0.05$ in t-test). Adding CSERP data not only fine-tunes the alignment of dialogue and profile across these five dimensions but also enhances the model’s logical and contextual coherence abilities through reasoning training, which is well reflected in the Coherence metric.
Overall, these results emphasize the importance of aligned dialogues and role profiles in role-playing training and demonstrate the excellent performance of the “Beyond Dialogue” training tasks in general role-playing dimensions.
Further baseline results can be found in the supplementary tables provided in §B. Case studies are available in §G. For a detailed account of the manual verification process used in GPT-4o evaluations, please refer to §C.
Alignment Results
To explore the model’s capability in fine-grained alignment for dialogue and role profiles, we randomly sampled 100 data points from each dimension of CSERP in the evaluation results (§Main Results). Using GPT-4o’s results as the reference labels, we present the alignment results of the comparable baseline model and two improved models in Table 3. The alignment for Character, Style, and Personality was assessed using recall, while Emotion and Relation were evaluated using NMAPE.
Model Character ↑ Style ↑ Emotion ↓ Relationship ↓ Personality ↑ Avg. ↑
Comparable Baselines
Yi-1.5-9B-Chat 79.0 ± 2.3 75.4 ± 2.5 17.1 ± 0.9 15.3 ± 1.5 70.7 ± 2.4 78.5 ± 1.9
GLM-4-9-chat 72.2 ± 2.3 77.8 ± 2.2 12.6 ± 0.7 12.7 ± 1.2 78.0 ± 2.4 80.5 ± 1.7
Custom Trained Baselines
Mistral-Nemo-Instruct-2407 52.8 ± 3.6 54.3 ± 4.2 13.5 ± 0.7 10.9 ± 0.8 77.2 ± 2.3 71.9 ± 2.3
+ RPA & CC & CSERP 89.1 ± 1.6 88.7 ± 2.2 10.1 ± 0.5 7.4 ± 0.8 87.0 ± 1.7 89.4 ± 1.4
[…]
+ RPA & CC & CSERP 86.3 ± 2.1 81.9 ± 2.5 13.1 ± 0.7 8.9 ± 0.9 82.7 ± 1.8 85.7 ± 1.6
Table 3: Comparison of Baseline Models on Dialogue and Role Profile Alignment Tasks. The results are benchmarked against GPT-4o’s alignment performance.
The fine-tuned Mistral-Nemo model, using the full dataset, achieved top performance, closely matching GPT-4o’s 89.4% effectiveness. Both Mistral and Qwen showed notable improvement, with recall rates for Character and Style increasing by over 30%, while NMAPE for Emotion and Relation decreased significantly. This enhancement allowed the 7B-parameter Qwen2 model to surpass the 9B-parameter Yi1.5 and GLM4 models in alignment tasks.
However, when trained on the RPA & CC or RP & CC datasets alone, both Qwen and Mistral struggled with alignment tasks. Their reasoning capabilities declined, resulting in ineffective outcomes. Detailed case studies are in §G.
The results from Table 2 and Table 3 indicate that enhancing the model’s ability to align dialogue and role profiles effectively translates into improved role-playing capabilities. Since these Alignment Tasks are consistent with the Evaluation Tasks across the five dimensions of CSERP, improvements in alignment ability also signify enhanced model role-playing evaluation capabilities. This is beneficial for establishing an effective feedback-enhanced loop for the role-playing.
Ablation Study
To assess the contribution of the five CSERP alignment tasks within the Beyond Dialogue framework, we conducted an ablation study. Each task was removed individually, with its training data volume replaced by an equivalent amount of CC data to maintain consistency. Subsequently, we evaluated the model across the five role-playing dimensions of CSERP, using the same evaluation methods as in Table 2.
Table 4 presents the results of the ablation experiments for CSERP training tasks. Notably, the ablation of w/o Char. and w/o Styl. have the most significant impact, with a marked decline in the performance of Character and Style, respectively. The other three metrics also showed varying degrees of decline, indicating that these fine-grained alignment tasks are crucial for the model’s general role-playing capability.
(a) Average Score
(b) Ablation Dimension
Figure 4: Comparison of Scores for Ablation Across Five Dimensions in CSERP (Left) and Performance Scores for Specific Dimensions (Right).
Model Character ↑ Style ↑ Emotion ↓ Relationship ↓ Personality ↑ Avg. ↑
w/o Char. 73.75 ± 1.24 81.43 ± 1.40 15.94 ± 0.45 11.75 ± 0.66 67.08 ± 4.56 78.91 ± 1.66
w/o Style. 76.94 ± 1.22 78.99 ± 1.38 15.87 ± 0.45 12.07 ± 0.69 67.25 ± 4.42 79.05 ± 1.63
w/o Emot. 78.69 ± 1.22 80.11 ± 1.38 16.08 ± 0.46 11.82 ± 0.65 67.58 ± 4.92 79.70 ± 1.73
w/o Rela. 77.69 ± 1.29 82.33 ± 1.39 15.81 ± 0.46 12.28 ± 0.68 68.92 ± 4.48 80.17 ± 1.66
w/o Pers. 78.38 ± 1.20 79.34 ± 1.42 16.14 ± 0.45 11.93 ± 0.66 66.17 ± 4.49 79.16 ± 1.64
All 78.67 ± 1.12 82.52 ± 1.33 15.68 ± 0.50 11.22 ± 0.72 69.67 ± 4.27 80.79 ± 1.59
Table 4: Ablation Results on CSERP Training Tasks.
Figure 4 presents a comparison of scores across the five dimensions of the CSERP model, illustrating how the ablation of individual tasks impacts both the overall performance and the specific dimension performances.
The ablation study clearly demonstrates that each alignment task within the Beyond Dialogue framework uniquely contributes to the model’s overall performance. The integration of all tasks yields the best results, underscoring the importance of these alignment tasks in training a robust general role-playing model.
Conclusion
This paper introduces a novel general role-playing framework called Beyond Dialogue. We propose a straightforward method that effectively aligns dialogues in specific scenarios with role profiles. This alignment helps to eliminate biases arising from inconsistencies between predefined profiles and the dialogues generated during training. Additionally, we present an innovative prompting mechanism that constructs “Beyond Dialogue” training tasks by generating reasoning processes. It enables fine-grained alignment between role profiles and dialogues at the sentence level. Experimental results demonstrate that our approach enhances the model’s ability to follow predefined profiles across various dimensions of general role-playing, surpassing most general and specialized role-playing baselines. We believe that our framework will inspire further exploration of diverse “Beyond Dialogue” training tasks in the field of general role-playing, thereby enhancing the ability of LLMs to effectively customize proprietary role models.
Appendix A Details of Dataset Construction
Dataset construction process
We need to manually collect and annotate character traits, speaking styles, and personality labels from the novels. The personality labels will be primarily sourced from the Personality Database [^1] https://www.personality-database.com/. Following this, we will proceed with the fully automated role-playing dialogue data construction process (see Figure 3):
  1. 1.
    Chunk Split and Filter: To manage the token limit of LLMs, the novel is split into fixed-token chunks. Only chunks that meet the role’s appearance threshold are retained, reducing the number of chunks and associated costs in subsequent stages, while ensuring that relevant roles are captured.
  2. 2.
    Scenario Extract and Chunk Evaluation: Open-source large models are utilized to extract scenarios from the chunks (prompt in Table 10), filtering out those that contain multiple scenarios to avoid discontinuous dialogues. Given the large number of input chunks, using specialized models can be costly. After comparison, we selected Qwen1.5-72B-Chat (Bai, Bai, and et al. 2023) for its balance of performance and efficiency. Additionally, chunks are evaluated to keep only those that reflect the role’s profile (prompt in Table 11), helping to eliminate non-matching chunks and reduce the number of chunks in subsequent inputs.
  3. 3.
    Dialogue Extract: Utilizing powerful LLMs to extract dialogues and actions between roles from valid chunks is crucial (prompt in Table 12), as role-playing dialogue data is fundamental for both role-playing training and subsequent derivative tasks. After comparing mainstream models, we selected GPT-4o (gpt-4o-2024-05-13) for its cost-effectiveness. §A details the extraction accuracy and cost of each LLM we evaluated.
  4. 4.
    Dialogue Check: To ensure valid role-playing dialogue data, we first focus on forming alternating two-person dialogues, which often leads to scenarios shrinking. This necessitates simultaneous scenario reconstruction and coherence checks to maintain contextual integrity (prompt in Table 13). After ensuring the coherence, we employ GPT-4 to review role portrayal accuracy, discarding any dialogues that conflict with the established profiles (prompt in Table 14).
  5. 5.
    Profile Alignment: We use GPT-4o to analyze each multi-turn dialogue, ensuring alignment of the role-playing data across the five dimensions: Character, Style, Emotion, Relationship, and Personality (CSERP, see Figure 5):
    1. (a)
      Character and Style: These are aligned using a word recall method, where GPT-4o recalls descriptive words from the role profile that are reflected in the dialogue within the scenario’s scope (prompt in Table 15 and Table 16).
    2. (b)
      Emotion and Relationship: These are aligned using a scale method, with GPT-4o providing a score from 0-10 based on the dialogue content. The Emotion dimension is based on Ekman’s six basic emotions (Ekman 1992): Anger, Disgust, Fear, Happiness, Sadness, and Surprise (prompt in Table 17). The Relationship dimension reflects the intimacy presented in the dialogues within the scenario’s scope (prompt in Table 18).
    3. (c)
      Personality: We use the Myers-Briggs Type Indicator (MBTI) [^2] https://www.16personalities.com/, which consists of four dimensions, each with two types. GPT-4o performs binary classification alignment for the four dimensions within the scenario’s scope (prompt in Table 19).
  6. 6.
    Profile Adjustment: In the role profile, Character, Style, and Personality are predefined, so we adjust each dialogue’s profile prompt based on the alignment results from Step 5. Since Emotion and Relationship are scene-dependent and cannot be predefined, we need to supplement the profile prompting the dialogue with information on these two dimensions.
Figure 5: An diagram of Deriving the Character-Style-Emotion-Relationship-Personality (CSERP) Dataset from Role-Playing Dialogue Data.
Comparative Analysis of Dialogue Extraction
This section presents a comparative analysis of LLMs for novel dialogue extraction, focusing on accuracy and cost. The models compared include Qwen1.5-72B-Chat (Bai, Bai, and et al. 2023), ERNIE-Bot (Baidu 2024), deepseek-chat (Deepseek 2024), moonshot-v1-8k (Moonshot 2024), and GPT-4o (OpenAI 2024). Key metrics are dialogue extraction accuracy, omission recall rate, and API costs. Note that Qwen, being open-source, incurs no API costs. A detailed comparison is provided in Table 5.
model Recall Accuracy Pricing (\$)
Qwen1.5-72B-Chat 67.7% 71.4% -
ERNIE-Bot 85.4% 91% 1.66 / 1M input tokens
1.66 / 1M output tokens
deepseek-chat 83.2% 92.2% 0.14 / 1M input tokens
0.28 / 1M output tokens
moonshot-v1-8k 73.1% 79.2% 1.66 / 1M input tokens
1.66 / 1M output tokens
gpt-4o 89.1% 96.4% 5.00 / 1M input tokens
15.00 / 1M output tokens
Table 5: Comparative Accuracy and Cost Analysis of Dialogue Extraction LLMs
High accuracy and recall rates are essential for maintaining the integrity and consistency of extracted dialogues. The quality of role-playing dialogue data significantly affects the quality of derived CSERP tasks. Compared to GPT-4o, the accuracy gap in the other LLMs is substantial, resulting in a notable decrease in the proportion of extracted dialogues passing the quality check. Consequently, producing the same amount of high-quality role-playing dialogue data with these LLMs incurs higher costs. Therefore, we ultimately chose GPT-4o for extracting character dialogues and actions from text segments.
Statistical Analysis of Beyond Dialogue Role-Playing Dataset
Following the method outlined in §A, we extracted 280 Chinese roles and 31 English roles from 123 Chinese and English novels or scripts. In total, 3,552 sessions of scenario dialogues were obtained, comprising 23,247 dialogue turns, all drawn from authentic dialogues in novels or scripts (Table 6).
Alphabetical List of Novel and Script Titles in English
A Dream of Splendor, A Record of a Mortal’s Journey to Immortality, A Slight Smile is Very Charming, A Song of Ice and Fire, All is Well, Battle Through the Heavens, Better Days, Better Days (Novel), Big Shot, Black Moonlight Hold Firm (Drama Script), Blade of the Immortal, Border Town, Bright Sword, Butterfly, Can’t Hide Love, Candle in the Tomb, Chinese Paladin 3, Chronicles of a Blood Merchant, Crouching Tiger, Day and Night, Deep Love and Rainy Weather, Demi-Gods and Semi-Devils, Detective Chinatown 2, Diamond Lover, Do You Know? The Green Should be Plump and the Red Lean, Dream of the Red Chamber, Empresses in the Palace, Ever Night, Farewell My Concubine, Fighter of the Destiny, First Love, Grandmaster of Demonic Cultivation, Guo Degang’s Comedy Collection, Half-demon Tsukasa, Hamlet, Handsome Siblings, Hard to Coax, Harry Potter, Heaven Official’s Blessing, Hello Mr. Billionaire, Hi, Hidden Dragon, Home with Kids (Season 1 Episodes 1-2), How Long Will I Love U, Howl’s Moving Castle, IPartment 1, IPartment 2 (Drama Script Excerpt), Important Things in Life, In the Mood for Love, In the Name of the People, Infernal Affairs, Journey Under the Midnight Sun, Journey to the West, Joy of Life (Novel), Joy of Life (Season 1 Episodes 1-2), Kong Yiji, Kung Fu, Longing Heart, Love As The Goal, Love in a Puff, Meeting You, Meteor, Mom (Drama Script Excerpt), My Homeland, My Own Swordsman, My People, Nirvana in Fire, Nirvana in Fire, Not Allowed to Die, Ode to Joy (Drama Script Excerpt), One and Only, Outlaws of the Marsh, Passing by Your World, Proud Wanderer, Qin’s Moon, Red Sorghum Clan, Silent Separation, Snow in Midsummer, Soldier Assault, Soul Land, Spirited Away, Sword, The Adventures of Chu Liuxiang, The Bad Kids, The Bride with White Hair, The Deer and the Cauldron, The Devotion of Suspect X, The Flowers of War, The Four Great Constables, The Godfather, The Great Dao Commander, The Heaven Sword and Dragon Saber, The Journey of Flower, The King’s Avatar, The Left Ear, The Legend of Lu Xiaofeng, The Legend of the Condor Heroes, The Liancheng Swordsman, The Longest Day in Chang’an, The Lost Tomb, The Lotus Flower Pavilion, The Masterless Master, The Mute Mansion, The Mystic Nine, The Newsroom (Drama Script Excerpt), The Orchid’s Oath, The Parasitic Son-in-Law, The Return of the Condor Heroes, The Romance of Tiger and Rose, The Smiling, The Speed of Life, The Story of Yanxi Palace , The Story of the Cook’s Camp (Drama Script Excerpt), The Sword Stained with Royal Blood, The Sword of the Third Young Master, The Swordsman, The Three-Body Problem, The Wandering Earth, The World Beneath (Text Excerpt), Tiny Times, To Live, To Try the World, To the Place with Wind (Drama Script Excerpt), Unrequited Love, White Deer Plain, With the Family (Drama Script Excerpt), You Are My Glory, Youth, Zhu Xian
Table 6: List of Novels and Scripts Titles
Table 7 summarizes the statistical information of various role-playing dialogue datasets. From the table, it can be seen that our dataset is entirely sourced from novels. However, this source provides richer and more authentic dialogue scenarios. Additionally, compared to other datasets, we have the highest number of real roles and the most sessions of authentic dialogues.
Dataset Source Open-source Multi-lingual Multi-turn # Roles # Sessions # Turns # Ave Turns
HPD Novel - 1042 14380 13.8
CharacterGLM Novel & Human & GPT 250 1034 16316 15.78
RoleLLM GPT Zh: 5, En: 95 - 23463 -
CharacterLLM GPT 9 1600 21120 13.2
PIPPA Human 1254 26000 1049015 40.34
ChatHaruhi Novel & GPT 32 54726 67660 1.23
WIKIROLE GPT Zh: 3184, En: 3902 7086 36164 5.1
Ours Novel Zh: 280, En: 31 3552 23247 6.54
Table 7: Comparison of Role-Playing Dialogue Datasets: Our dataset vs. Existing Role-Playing Datasets. Note: In the HPD dataset, the number of roles is denoted as “-” since it is exclusively centered on Harry Potter, with other characters interacting with him. RoleLLM is single-turn, so # Sessions and # Avg Turns are marked as “-”.
In Figure 6, the distribution of dialogue turns in both Chinese and English from our role-playing dataset is presented, illustrating the variation in dialogue lengths.
(a) Chinese
(b) English
Figure 6: Distribution of Dialogue Turns in Our Role-Playing Dialogues Dataset
The verb-noun structure of role-playing instructions in English is illustrated in Figure 7, with the inner circle representing the top 20 verbs and the outer circle listing the direct noun objects.
Figure 7: Verb-noun Structure of Our Role-playing Instructions in English.
The word clouds in Figure 8 illustrate the character traits and speaking styles present in our role profiles.
(a) Character
(b) Styles
Figure 8: Word Cloud of Character Traits and Speaking Styles in Role Profiles
As illustrated in Figure 9(a) and 9(b), the distribution of emotion and relationship values is presented, both rated on a scale from 0 to 10. The emotion distribution is based on the highest emotion score for each dialogue. Additionally, Figure 9(c) shows the varied distribution of MBTI personality types among the role profiles in the dataset.
(a) Emotion
(b) Relationship
(c) Personality
Figure 9: Distribution of Emotion Values, Relationship Values, and MBTI Personality Types Among Role Profiles in the Dataset
Appendix B Supplementary Experimental Results
Supplementary Baselines for Main Results
This appendix complements the main results presented in the paper by providing additional insights into the bilingual chatbot baselines used in our evaluation. In addition to our proprietary general baselines, we have also tested newly added open-source general baselines. The proprietary models include GPT-4o, GPT-3.5-Turbo (OpenAI 2024), Moonshot-v1-8k (Moonshot 2024), Yi-Large-Turbo (Wanwu 2024), Deepseek-Chat (Deepseek 2024), Baichuan4 (Baichuan 2024), and Hunyuan (Tencent 2024). The newly tested open-source baselines include Yi-1.5-9B-Chat, GLM-4-9b-chat, Mistral-Nemo-Instruct-2407, and Qwen2-7B-Instruct.
Based on the experimental results, we selected Qwen2-7B-Instruct and Mistral-Nemo-Instruct-2407 for our training due to their balanced performance across both the average score (”Avg.”) and human likeness dimensions. While Yi-1.5-9B-Chat demonstrated a reasonably high average score, its performance in human likeness was notably poor. Its dialogues were excessively rigid and AI-like, which compromised the immersion and authenticity needed in role-playing scenarios.
Moreover, an important factor in selecting Mistral-Nemo-Instruct-2407 was its primary focus on English, in contrast to the other three models—Yi-1.5-9B-Chat, GLM-4-9b-chat, and Qwen2-7B-Instruct—which are predominantly Chinese-language models. Including Mistral in our training set was crucial for ensuring linguistic diversity, thereby enhancing the model’s capability to handle bilingual tasks effectively.
Model Character Style Emotion Relationship Personality Avg. ↑ Human -likeness ↑ Role Choice ↑ Coherence ↑
Recall ↑ Recall ↑ NMAPE ↓ NMAPE ↓ Precision ↑
General Baselines(Proprietary)
GPT-4o 74.32 ± 1.15 81.67 ± 1.51 16.31 ± 0.48 12.13 ± 0.66 66.58 ± 4.41 78.83 ± 1.64 67.33 ± 3.95 87.33 ± 3.86 99.67 ± 0.33
GPT-3.5-Turbo 72.26 ± 1.27 73.66 ± 1.73 17.79 ± 0.56 14.17 ± 0.73 66.92 ± 4.85 76.18 ± 1.83 $3$ 4.43 83.00 ± 4.68 97.33 ± 1.17
Moonshot-v1-8k 74.06 ± 1.19 80.64 ± 1.51 16.17 ± 0.47 13.42 ± 0.70 67.00 ± 4.87 78.42 ± 1.75 44.00 ± 4.33 86.67 ± 3.75 99.33 ± 0.46
Yi-Large-Turbo 75.13 ± 1.22 79.18 ± 1.58 16.44 ± 0.49 13.48 ± 0.67 68.25 ± 4.61 78.53 ± 1.72 47.00 ± 4.60 84.33 ± 3.67 92.67 ± 2.39
Deepseek-Chat 75.46 ± 1.14 81.49 ± 1.51 15.92 ± 0.46 12.42 ± 0.63 67.92 ± 4.57 79.30 ± 1.66 52.33 ± 4.95 83.00 ± 4.68 96.67 ± 1.00
Baichuan4 71.82 ± 1.25 76.92 ± 1.52 17.57 ± 0.52 12.30 ± 0.62 67.08 ± 4.75 77.19 ± 1.73 45.33 ± 4.31 82.33 ± 4.49 99.33 ± 0.46
Hunyuan 73.77 ± 1.18 78.75 ± 1.56 17.24 ± 0.48 13.22 ± 0.68 67.00 ± 4.39 77.81 ± 1.66 53.00 ± 4.29 84.33 ± 4.52 98.33 ± 0.84
General Baselines(Open-source)
Yi-1.5-9B-Chat 75.31 ± 1.20 76.78 ± 1.49 16.67 ± 0.52 12.75 ± 0.66 67.42 ± 4.63 78.02 ± 1.70 38.67 ± 4.39 84.00 ± 4.61 92.67 ± 1.79
GLM-4-9b-chat 74.26 ± 1.19 78.40 ± 1.55 17.18 ± 0.50 14.48 ± 0.74 67.17 ± 4.93 77.63 ± 1.78 47.67 ± 4.25 83.33 ± 4.51 99.33 ± 0.46
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 77.53 ± 1.61 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
Qwen2-7B-Instruct 75.39 ± 1.13 77.68 ± 1.65 17.64 ± 0.56 13.43 ± 0.7 67.75 ± 4.44 77.95 ± 1.70 48.00 ± 4.66 83.33 ± 4.48 99.00 ± 0.56
Table 8: Supplementary Main Results of Beyond Dialogue. Bold values indicate the best performance in each group.
Supplementary Result for Ablation Study
Figure 10: Radar Chart of Ablation Results on CSERP Training Tasks.
The radar charts in Figure 10 visually encapsulate the comprehensive impact of each training task’s absence, offering a nuanced view of their contributions to the model’s overall capability across multiple dimensions. By systematically removing individual CSERP tasks and observing the resulting changes in performance, this analysis highlights the critical role each task plays in enhancing the model’s alignment with the role-playing dimensions. The significant drops in Character and Style metrics upon the removal of w/o Char. and w/o Styl. tasks, respectively, emphasize the necessity of these fine-grained alignment tasks in maintaining the model’s general role-playing effectiveness.
Appendix C Human Verification of Evaluation
In this section, we present our approach to verifying the evaluations conducted by GPT-4o through human assessment. Specifically, we sampled 100 dialogue evaluations generated by GPT-4o and manually checked the reasonableness of these evaluations. The dialogue evaluations were divided into two parts: Auto-Generation and Auto-Evaluation.
For the Auto-Generation part, GPT-4o’s task was to generate dialogue scenarios and corresponding scenario settings. Human judges evaluated the validity of the content generated by GPT-4o. The validity metric reflects the percentage of instances deemed reasonable and accurate by human evaluators. As shown in Table 9, the validity scores across different metrics such as Role, Scenario, Emotion, Relationship, and Dialogue are consistently high, indicating that GPT-4o performs well in generating contextually appropriate dialogue scenarios.
In the Auto-Evaluation part, GPT-4o evaluated role-playing dialogues generated by different models across eight dimensions: Character, Style, Emotion, Relationship, Personality, Human-likeness, Role Choice, and Coherence. The human verification process involved two key metrics: Validity and Cosine similarity. Validity assesses the soundness of GPT-4o’s evaluation justifications, while Cosine similarity measures the agreement between GPT-4o’s evaluation results and those of human evaluators.
As depicted in Table 9, the validity scores for Auto-Test dimensions are consistently high, with the lowest score being 0.88, which reflects the robustness of GPT-4o’s evaluation justifications. However, the Cosine similarity metric indicates that while there is strong alignment between GPT-4o’s evaluations and human assessments in dimensions such as Relationship, Personality, and Coherence, there is a notable discrepancy in the Human-likeness dimension, with a Cosine similarity of only 0.78. This suggests that GPT-4o’s evaluations of human likeness deviate more significantly from human evaluations compared to other dimensions. Additionally, the evaluation of human likeness is likely influenced by biases in individual human judgments, contributing to this discrepancy.
Auto-Generation Auto-Evaluation
Metrics Role Scenario Emotion Relationship Dialogue Character Style Emotion Relationship Personality Human-likeness Role Choice Coherence
Validity 1.0 ± 0.0 1.0 ± 0.0 0.94 ± 0.02 1.0 ± 0.0 1.0 ± 0.0 0.89 ± 0.04 0.9 ± 0.03 0.92 ± 0.03 0.98 ± 0.01 0.93 ± 0.03 0.88 ± 0.03 0.92 ± 0.03 0.99 ± 0.01
Cosine - 0.96 ± 0.01 0.94 ± 0.01 0.94 ± 0.01 0.99 ± 0.0 0.98 ± 0.01 0.78 ± 0.0 0.91 ± 0.03 0.99 ± 0.0
Table 9: Human Verification of GPT-4o’s Evaluation in Auto-Generation and Auto-Test. Validity reflects the percentage of evaluations deemed reasonable by human evaluators. Cosine similarity measures the agreement between GPT-4o’s evaluation results and those of human evaluators.
The results demonstrate that GPT-4o’s evaluations are largely consistent with human judgments, showcasing its capability to perform reliable and contextually appropriate assessments in role-playing dialogue scenarios.
Appendix D Implementation Details
Training. We fully finetuning all models for 3 epochs with 833 steps with the 128 global batch size. We set the training sequence length to 4,096. The learning rate is 3e-5. The + RPA & CC & CSERP represents models were trained using Align Role-Playing dialogue with profiles, Chit-Chat, and CSERP data in a ratio of 1:4:5, while + RP & CC represents a ratio of 1:9 using non-aligned dialogue data. In both cases, the proportion of English to Chinese data within the Chit-Chat data is consistent with that in the Role-Playing data.
Inference. We infer all models with topP 0.8, length penalty 1.1, sequence length 4096, and max new token 256. The number of dialogue turns in the automated dialogue setting is 5. During the inference of the Alignment task, some outputs may fail to parse due to formatting requirements. We allow the model five inference attempts. Any result that meets the formatting criteria at least once is retained to evaluate its alignment capabilities in dialogue and role profile.
Evaluation. We evaluated a total of 30 roles, with 15 roles drawn from the training set and 15 newly defined roles. Each role engaged in dialogues with 10 randomly generated new roles, with each dialogue consisting of 5 turns. The ratio of Chinese to English roles in the evaluation was 2:1, consistent with the ratio in the training data.
Hardware and System Configuration We use 8 NVIDIA RTX A100 GPUs with 1TB of memory for training and a single A100 GPU for inference. The system operates on Linux version 4.14.105-1-tlinux3-0013.
Appendix E Cost Analysis
In this section, we provide a detailed cost analysis of the Beyond Dialogue framework, including both automated and manual evaluation processes. For the manual evaluation, we enlisted the assistance of three evaluators from Mainland China, all of whom possess proficient English skills, ensuring the accuracy and reliability of the assessments. The cost of human annotation was calculated at \$16/hour (approximately \$1 per data item), a rate that exceeds the minimum wage in Mainland China. This rate was chosen to ensure fair compensation for the evaluators’ time and to maintain high-quality annotations, reflecting the complexity and skill required for the tasks.
Role-playing Dialogue Dataset Construction
In this section, we calculated the costs associated with constructing the role-playing dialogue dataset, ensuring both cost-effectiveness and efficiency.
Automated Pipeline. We employed GPT-4o for dialogue extraction and reconstruction, with an average cost of \$0.016 per item. Initial chunk checking and scenario construction were handled by open-source models, and excluded from the cost analysis. The comprehensive extraction and verification process using GPT-4o amounted to \$0.048 per item.
Human Annotation. For comparison, human annotation costs were \$16/hour (approximately \$1 per item), significantly higher than the automated method. Human annotation not only incurs over 20 times the cost per item but also is substantially slower.
Our findings highlight that the pipeline we employed for constructing the role-playing dialogue dataset is both low-cost and highly efficient. Despite the significantly lower cost, we ensured the reliability of the automated process by conducting a thorough human review of the dialogues extracted by GPT-4o. As detailed in §A, the accuracy of GPT-4o’s outputs was found to exceed 96%, demonstrating that the constructed dataset is trustworthy and of high quality. In conclusion, our automated pipeline not only reduces costs but also accelerates the process of dataset construction, providing a scalable solution for future projects while maintaining a high standard of data accuracy.
Alignment and Evaluation
In this section, we present a detailed cost analysis of the alignment and evaluation process for Beyond Dialogue framework. Leveraging our automated pipeline, we conducted a comprehensive evaluation across multiple dimensions, including Character, Style, Emotion, Relationship, and Personality, as well as additional tasks such as Human-likeness, Coherence, and Role Choice. The key components of our cost analysis are outlined below.
Automated Alignment. The alignment of data across these five dimensions (Character, Style, Emotion, Relationship, and Personality) was performed using our automated pipeline, with an average cost of \$0.012 per data item. In contrast, manual annotation for alignment would cost approximately \$16 per hour, which translates to about \$0.5 per data item. This demonstrates that our automated method reduces the cost by a factor of over 40 compared to traditional human annotation.
Automated Evaluation. Before the evaluation, our automated process includes an initial cost for generating dialogues with the model. Each model interaction involves a 5-round dialogue, costing approximately \$0.02 per data item. Following this, our evaluation process expands beyond the alignment dimensions to include three additional task(Human-likeness, Coherence, and Role Choice), which are essential for ensuring that the dialogues not only align with predefined profiles but also exhibit realism and logical consistency. The cost of evaluating each data item across these eight dimensions, including the initial model interaction, is approximately \$0.1 per data item. For comparison, manual evaluation, which involves interacting with the model and performing the eight evaluation tasks, takes about 12 minutes per data item, resulting in a cost of approximately \$5 per item. This comparison highlights that our automated evaluation process is over 50 times more cost-effective than traditional human annotation.
This evaluation process, which includes both automated alignment and additional assessment dimensions, underscores the efficiency and cost-effectiveness of our pipeline. By leveraging automation, we achieve a comprehensive evaluation at a fraction of the cost and time required for manual annotation, while still maintaining high standards of accuracy and quality, as demonstrated in §C. In conclusion, our alignment and evaluation framework not only ensures that the constructed role-playing dialogue dataset is aligned with the intended profiles but also that it is evaluated rigorously for quality and coherence, all while significantly reducing costs compared to traditional methods.
Appendix F Prompt Templates
Dataset Construction Prompts
In the process of constructing the role-playing dialogue dataset, a variety of prompt templates were strategically utilized to extract and evaluate relevant scenarios and dialogues, ensuring both the accuracy and coherence of the dataset. By minimizing human intervention, the process remains highly low-cost. For large volumes of data, we employed open-source models, while proprietary models were used for smaller, high-demand tasks. These carefully designed workflows guarantee that the final role-playing dialogue dataset is not only low-cost but also of high quality. Below are the key prompt templates used:
  • Prompt Template for Extracting Chunk Scenarios: This template, as shown in Table 10, is designed to identify and extract meaningful scenarios from chunks of text.
  • Prompt Template for Evaluating Chunk Based on Character Traits: As illustrated in Table 11, this template is used to assess the relevance and quality of the extracted chunks with a focus on character traits.
  • Prompt Template for Extracting Dialogues: Table 12 presents the template utilized for identifying and extracting dialogues from the chunks.
  • Prompt Template for Scenario Reconstruction and Coherence Checking: To ensure the logical consistency of the scenarios, this template, shown in Table 13, is employed for reconstructing scenarios and checking coherence.
  • Prompt Template for Dialogue Conflict Detection: Finally, Table 14 highlights the template used for detecting conflicts within dialogues, ensuring that the dataset remains coherent and free of contradictions.
Alignment Prompts
In the construction of role-playing training data, a special emphasis is placed on accurately aligning the personality profile (prompt) with the dialogue content (label) across multiple dimensions. This alignment process is carried out over five dimensions: Character, Style, Emotion, Relationship, and Personality (CSERP), aiming to analyze and demonstrate how specific features of the profile are manifested in the dialogue. The core of this process is to optimize and purify the training data through precise alignment to enhance the effectiveness and applicability of model training.
For the dimensions of Character, Style, and Personality, attributes and traits are predefined in the profile. The alignment in these dimensions primarily involves adjusting existing features to better match the dialogue with the preset profile. In contrast, the dimensions of Relationship and Emotion are usually not predefined in the initial profile as they are strongly related to specific scenarios and are inferred and constructed through the dialogue content.
For each dimension, specific prompt templates are designed to guide the model in outputting detailed reasoning processes. For example, the model needs to identify and explain which sentence in the dialogue displays which trait from the profile, thus achieving sentence-level alignment between the profile and the dialogue.
The prompt templates for each dimension are detailed as follows:
  1. 1.
    Character Alignment: This prompt focuses on analyzing and identifying character traits from dialogue content in relation to predefined character candidates (Table 15).
  2. 2.
    Style Alignment: This prompt analyzes a character’s speaking style from dialogue content to match it with specific style candidates (Table 16).
  3. 3.
    Emotion Alignment: This prompt requires the analysis of Ekman basic emotions (Ekman 1992) from the dialogue of a specified role, assessing emotional expressions in context to the scene (Table 17).
  4. 4.
    Relationship Alignment: This prompt evaluates character intimacy through dialogue and scene interactions (Table 18).
  5. 5.
    Personality Alignment: This prompt focuses on determining the MBTI personality type of a character by analyzing dialogue content and role information (Table 19).
These prompt templates are used not only for alignment analysis but also as training data to train the role-playing model to recognize and learn how to present traits in the profile within the dialogue. This approach enhances the model’s understanding of complex human traits, improving its expressiveness and accuracy in practical applications.
Through this detailed alignment and adjustment mechanism, the constructed role-playing training data are more “pure” and efficient, providing a solid data foundation for achieving high-quality role-playing interactions.
Auto Dialogue and Evaluation Prompts
The prompt templates used in the automated dialogue pipeline are as follows: chat role generation (Table 20), dialogue scenario generation (Table 21), and generation of Emotion and Relationship (Tables 22 and 23). Additionally, prompts for playing the chat role and prompts for playing the evaluated roles are presented in Tables 24 and 25.
In our role-playing training data framework, evaluation prompts aim to assess dialogues for adherence to established profiles, reflecting the quality of role portrayal. To this end, evaluation prompts for the CSERP dimensions (Character, Style, Emotion, Relationship, Personality) align with the templates in Alignment Prompts in §F.
Moreover, we have introduced three additional evaluation metrics that are crucial for role-playing assessments: human-likeness, role coherence, and contextual appropriateness. Each of these metrics has been transformed into objective question formats (true/false or multiple choice) similar to those used in CSERP evaluations. The evaluation prompts for these metrics are detailed as follows:
  1. 1.
    Prompt for Human-likeness Evaluation (Table 26): This prompt assesses whether the dialogue samples resemble authentic human interaction. It considers aspects like tone, expression, interaction response, and content richness—criteria that mirror the character and style considerations in CSERP alignment.
  2. 2.
    Prompt for Role Choice Evaluation (Table 27): This prompt focuses on identifying the correct identity of dialogue participants based on their spoken content within a given scene. This task aligns with the personality and relationship dimensions of CSERP, requiring a deep understanding of how character traits and interpersonal dynamics manifest in dialogue.
  3. 3.
    Prompt for Coherence Evaluation (Table 28): This prompt examines the logical flow and contextual integration of dialogues within a scene. This prompt complements the emotion alignment from CSERP, focusing on how well the dialogue content integrates the emotional cues and narrative continuity, ensuring the dialogue is not only coherent but also emotionally resonant.
This standardization ensures that our assessments are not only efficient but also reproducible, allowing us to reliably measure and improve the fidelity and instructional value of our role-playing dialogues
Appendix G Case Study
In this section, we sample several cases from the role-playing dialogue and alignment tasks to illustrate the effectiveness of our proposed framework.
Role-playing Dialogue
In this subsection, we present a series of dialogue cases in both Chinese and English, with the original Chinese dialogues translated into English. The cases are illustrated in Figures 11 to 18, showcasing the role-playing capabilities of the Qwen2-7B and Mistral-Nemo models.
Through our case studies, we observe that the models, after training, exhibit significant improvements in adhering to the predefined role profiles and producing more human-like responses. For instance, the trained Mistral-Nemo model, when role-playing as Hamlet, consistently generates responses that are polite and eloquent, reflecting the role’s sophisticated nature. Similarly, the trained Qwen2-7B model, when portraying Zhou Botong in Chinese, captures the essence of his “Old Urchin” persona, demonstrating a mischievous and playful demeanor in its output. These results underscore the effectiveness of our training approach in enhancing the models’ ability to represent complex role traits in role-playing dialogues accurately.
Alignment Task
This subsection delves into the alignment task within the CSERP framework, which encompasses five critical dimensions: Character, Style, Emotion, Relationship, and Personality. For Character, Style, and Personality (CSP), alignment is assessed through word recall, while Emotion and Relationship (EP) are evaluated using a 0-10 scoring system. Thus, we selected one task from each of the aforementioned types for case demonstration, specifically focusing on Character and Emotion, as shown in Figures 19 to 22.
The untrained Qwen2-7B model struggles with fine-grained alignment of profiles and dialogues, often failing to follow instructions effectively. After training, both Mistral-Nemo and Qwen2-7B models perform closer to GPT-4o.
Prompt for Extracting Chunk Scenarios
You are an expert with a deep understanding of literary works, skilled at analyzing and extracting the core elements of literature. Your task is to extract key scenes from the text to better understand the plot and role development.
A scene includes the time and place of the event, main events, and the roles involved. Do not include role dialogues in the scene. Here is an example of a provided scene:
[Scene]
{scenario example}
Now we begin extracting scenes from a new text:
[Text]
{chunk}
[Requirements]
1. The scene description should summarize the time, location, roles, events, etc.
2. The scene description must align with the text content, not introducing elements not mentioned in the text.
3. The scene description should not include role dialogues.
4. The scene description should be between 100-150 words in length.
Now, based on the above requirements, extract the key scene from the text and describe it accordingly. Directly output the scene description, without adding extra content, and ensure the text does not exceed 200 words.
Table 10: Prompt Template for Extracting Chunk Scenarios.
Prompt for Evaluating Chunk Based on Character Traits
You are an expert with a strong background in literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances of roles from texts. Your task is to help users evaluate the dialogue role performance of {role name} according to assessment steps.
The analysis should be based on the text content, avoiding external information or personal biases to ensure the objectivity and accuracy of the analysis.
[Character Traits]
{character}
[Text]
{chunk}
[Evaluation Criteria]
Effectiveness (1-10): How well do the words spoken by {role name} in the text reflect {role name}’s character traits?
[Evaluation Steps]
1. Read and understand the role description.
2. Read and understand the text provided by the user.
3. Identify what the role has said in the text.
4. Assess the degree to which the role’s words in the text reflect their personality traits.
5. Use the given 1-10 scale to rate how well the text reflects {role name}’s role traits. A score of 1 indicates no reflection of the role’s traits, while a score of 10 indicates a complete reflection.
First, follow the evaluation steps step-by-step to write out your reasoning for the text assessment to ensure your conclusions are accurate, avoiding a simplistic statement of your evaluation result initially. Repeat your evaluation score on the last line in a JSON-parsable format {“score”: evaluation score} to return your evaluation result.
Table 11: Prompt Template for Evaluating Chunk Base on Character Traits.
Prompt for Extracting Dialogues
Your goal is to extract structured information from the user’s input that matches the form described below. When extracting information please make sure it matches the type information exactly. Do not add any attributes that do not appear in the schema shown below.
‘‘TypeScript
script: Array<{ // Adapted from the novel into script
role: string // The role who is speaking or performing an action, use context to predict the name of the role. Use scene or narrator if no one speak
dialogue: string // The dialogue spoken by the roles in the sentence, equals “-” if it’s no dialogue
action: string // The actions performed by the roles in the text, A high-level summary of a role’s behavior. action equals “dialogue”. if it’s no dialogue, summarize role’s behavior in sentence
Please output the extracted information in CSV format in Excel dialect. Please use a |||| as the delimiter.
Do NOT add any clarifying information. Output MUST follow the schema above. Do NOT add any additional columns that do not appear in the schema.
{extract example}
Input: {user input}
Output:
Table 12: Prompt Template for Extracting Dialogues.
Prompt for Scenario Reconstruction and Coherence Checking
You are an expert in scene analysis, skilled at analyzing and extracting key information from texts. Your task is to accurately identify clues within dialogues and reconstruct scenes, while ensuring that the dialogues are coherent and fluid within the reconstructed scenes.
- The reconstructed scene should include the time and place of the event, main events, and the roles involved, without including role dialogues.
- Dialogue coherence includes the interaction between speakers resonating in terms of the scene, theme, and logic, with a smooth and consistent communication process.
Here is a reference scene I provide. You need to identify the scene context of the provided dialogue and reconstruct its description:
[Scene]
{scene}
[Dialogue]
{dialogue}
Now, based on the above requirements, reconstruct the sub-scene where the dialogue takes place, and describe it accordingly. Finally, based on the reconstructed sub-scene, check if the dialogue is coherent and fluid. You may briefly analyze the scene context and its coherence, then return your evaluation result in a JSON-parsable format as follows:
{“scene”: “reconstructed scene description”, “coherence”: 1/0} Where “coherence” of 1 indicates the dialogue is coherent and fluid with the scene, and 0 indicates the dialogue is not coherent with the scene.
Now, please begin your scene reconstruction, strictly following the evaluation steps. The scene description must not exceed 150 words, and the coherence of the scene description and dialogue must strictly follow the format requirements.
Table 13: Prompt Template for Scenario Reconstruction and Coherence Checking.
Prompt for Dialogue Confilict Detection
You are an expert in the fields of literature and psychology, skilled at analyzing and interpreting the role traits and dialogue performances in texts. Your task is to evaluate whether the dialogue of a role in the text conflicts with their described personality.
[Role Description]
{role des}
[Scene]
{scene}
[Dialogue]
{dialogue}
[Evaluation Steps]
1. Read and understand the role description.
2. Read and comprehend the dialogue of the role.
3. Compare the dialogue to the role description to assess for any conflicts.
- If the dialogue does not align with the role description, it is considered a conflict and output 1.
- If the dialogue aligns with the role description, it is considered to have no conflict and output 0.
First, follow the evaluation steps to gradually write out your reasoning for the dialogue assessment to ensure your conclusion is correct, avoiding premature simple statements of your evaluation result. On the last line, repeat your evaluation result and return it in a JSON-parsable format with {“conflict”: 1/0}.
Table 14: Prompt Template for Dialogue Confilict Detection.
Prompt for Character Alignment
You are a character analysis expert, skilled in analyzing character traits from dialogue content and matching them to a provided set of character candidates.
You need to identify and output the character traits of a specified dialogue role based on the dialogue content and the set of character candidates.
[Scene]
{scene}
[Dialogues ]
{dialogues}
Based on the above dialogue content and scene, analyze the character traits of the {role name}. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
[Candidate Character Set]
{character candidates}
Return your evaluation result in a JSON-parsable format, with each character type separated by a comma. The specific format is as follows:
{”character”: “trait1, trait2…”}
Now, please begin your analysis of {role name}’s character. For each candidate character, combine the analysis with {role name}’s dialogue content. Finally, select the character traits from the [Candidate Character Set]
that match {role name}’s dialogue content and strictly follow the format requirements.
Table 15: Prompt Template for Character Alignment.
Prompt for Style Alignment
You are a professional speaking style analyst, skilled in analyzing characters’ speaking styles from dialogue content and matching them to a provided set of style candidates.
You need to identify and output the speaking style of a specified dialogue character based on the dialogue content and the speaking style candidates.
[Scene]
{scene}
[Dialogues]
{dialogues}
Based on the dialogue content and scene above, analyze the speaking style of the {role name}. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct.
[Candidate Speaking Styles]
{style candidates}
Return your evaluation result in a JSON-parsable format, with each speaking style separated by a comma. The specific format is as follows:
{“style”: “style1, style2…”}
Now, please begin your analysis of {role name}’s speaking style. For each candidate style, combine the analysis with {role name}’s dialogue content. Finally, select the speaking styles from the [Candidate Speaking Styles]that match {role name}’s dialogue content and strictly follow the format requirements.
Table 16: Prompt Template for Style Alignment.
Prompt for Emotion Alignment
You are an expert in the field of emotional psychology, skilled at analyzing emotions through a role’s dialogues, actions, and scenes.
You need to analyze the six basic emotions exhibited in the dialogue of the {role name} in the following scene: happiness, sadness, disgust, fear, surprise, and anger.
[Role Information]
{role name}’s character is {character}, MBTI type is {MBTI}, and speaking style is {style}.
[Scene]
{scene}
[Dialogues]
dialogues
Understand the role information and the current scene, and assess through the dialogues the degree to which {role name} exhibits the six basic emotions: happiness, sadness, disgust, fear, surprise, and anger in that scene. Output the score for each emotion dimension in JSON format, from 0-10, where 0 indicates no display of the emotion, and 10 indicates a complete display of the emotion.
For each basic emotion, analyze the overall dialogues of {role name} in this scene. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation results in a JSON-parsable format as follows:
{“happiness”: happiness score, “sadness”: sadness score, “disgust”: disgust score, “fear”: fear score, “surprise”: surprise score, “anger”: anger score}
Now, please begin your dialogue emotion analysis, and the final emotion scores must strictly follow the format requirements.
Table 17: Prompt Template for Emotion Alignment.
Prompt for Relationship Alignment
You are an emotional analysis expert, proficient in emotional analysis, psychology, dialogue understanding, and interpersonal relationship assessment. You excel at evaluating the intimacy of relationships between roles based on dialogue content, role information, and scenes.
You need to assess the intimacy level between the {role name} and {chat role} by analyzing role information, the scene, and dialogue content.
[Role Information]
{role name}’s character is {character}, MBTI type is {MBTI}, and speaking style is {style}.
[Scene]
{scene}
[Dialogues]
{dialogues}
Understand {role name}’s information, consider the current scene’s impact on role relationships, evaluate the overall dialogue content, focusing on the depth of emotional expression and interaction, and combine these factors to provide an intimacy score and analysis. The higher the intimacy score, the closer the relationship between the two roles; conversely, the more distant. The intimacy score ranges from 0-10, where 0 represents the most distant relationships, indicating strangers, hostility, indifference, etc., and 10 represents the closest relationships, such as lovers, kin, or friends.
Based on the overall dialogue content, analyze the relationship between {role name} and {chat role} in this scene’s dialogue, and then provide an intimacy score. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:
{“relationship”: intimacy score}
Now, please begin your intimacy assessment between {role name} and {chat role}, ensuring that the final intimacy score strictly follows the format requirements.
Table 18: Prompt Template for Relationship Alignment.
Prompt for Personality Alignment
You are an experienced psychologist skilled in analyzing role personalities through dialogue content and accurately determining MBTI personality types.
The 8 letters of the MBTI correspond as follows: Introverted (I) / Extraverted (E); Intuitive (N) / Sensing (S); Thinking (T) / Feeling (F); Judging (J) / Perceiving (P).
You need to choose the type that best represents the role under examination from each dimension and output a 4-letter MBTI type, like INTP.
[Role Information]
{role name}’s character is {character}, and speaking style is {style}.
[Scene]
{scene}
[Dialogues]
{dialogues}
Based on the above dialogues and scene, analyze the personality of the {role name} across the four MBTI dimensions. Ensure your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:
{“personality”: “MBTI type”}
Now, please begin your analysis of {role name}’s personality, and the final MBTI type must strictly follow the format requirements.
Table 19: Prompt Template for Personality Alignment.
Prompt for Chat Role Generation
You are an experienced creative writing tutor, skilled in creating innovative roles.
You need to design a new role description that will converse with {role name}, ensuring that the dialogue with this role effectively reflects {role name}’s personality, character traits, and speaking style.
Here is some basic information about {role name}:
Character: {character}
MBTI personality type: {MBTI}
Speaking style: {style}
World: {world}
You need to creatively construct a new role setting to dialogue with {role name}, based on the traits of {role name}. This new role should not appear in any works related to {role name}. The description of the new role should include the role’s name and a brief personal description, to be output in JSON format like:
{“chat role”: “role’s first name”, “role des”: “role’s description (not exceeding 100 words)”}
Below are some reference roles:
{reference}
Please design a completely new role that is distinctly different from these reference roles.
Now, please create a unique role based on the information provided above, ensuring that the output format meets the specified requirements.
Table 20: Prompt Template for Chat Role Generation.
Prompt for Scenario Generation
You are an experienced screenwriter skilled in creating engaging scenes.
You need to create a scene description that fits the settings of two roles while being consistent with the world in which the roles exist.
For reference:
- {scene example}
Here is some basic information about the dialogue role:
Role A:
Name: {role name}
Role description: {role name}’s character is {character}, MBTI personality type is {MBTI}, and speaking style is {style}.
Role B:
Name: {chat role}
Role description: {role des}
World of the roles: {world}
You need to construct an engaging scene based on the information of roles A and B. The scene and roles’ actions must be consistent with the settings of both roles, ideally within 50-100 words, and consistent with the world they inhabit. The output should be in JSON format, like
{“scene”: “scene description (50-100 words)”}
Now, please create a scene that fits the settings of both roles and is engaging. The scene should not directly include roles’ dialogues. Ensure that the output format meets the specified requirements.
Table 21: Prompt Template for Scenario Generation.
Prompt for Emotion Generation
You are a professional psychologist, skilled in analyzing role’s emotions and behavioral patterns.
You need to assign six basic emotions to the {role name} in a specific scene: happiness, sadness, disgust, fear, surprise, and anger, based on the dialogue role’s information and the scene of the dialogue.
Here is some information about the dialogue roles and the scene:
Role A:
Name: {role name}
Role description: {role name}’s character is {character}, MBTI personality type is {MBTI}, and speaking style is {style}.
Role B:
Name: {chat role}
Role description: {role des}
Scene:
{scene}
Understand the role descriptions and the current scene, and assign the six basic emotions reflected in {role name}’s statements in that scene: happiness, sadness, disgust, fear, surprise, and anger. Output the score for each emotion dimension in JSON format, from 0-10, where 0 means the emotion is not displayed at all, and 10 means the emotion is fully displayed.
Please analyze in a few brief sentences the scores for the six basic emotions that {role name} should exhibit in this scene, avoiding a simplistic statement of your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation results in a JSON-parsable format as follows:
{“happiness”: “happiness score”, “sadness”: “sadness score”, “disgust”: “disgust score”, “fear”: “fear score”, “surprise”: “surprise score”, “anger”: “anger score”}
Now, strictly follow the requirements to analyze the scene and role information, and assign emotional scores to {role name} that must be consistent with the role’s settings and the current scene. The analysis should be brief, not too lengthy, and avoid additional content. The final emotion scores must strictly follow the format requirements.
Table 22: Prompt Template for Emotion Generation.
Prompt for Relationship Generation
You are an emotional analysis expert, proficient in psychology and interpersonal relationship assessment, skilled at initializing the intimacy level of relationships based on roles’ personalities and scenarios. You need to analyze the intimacy level between the {role name} and {chat role} by analyzing the information of both roles and the content of the scene.
Here is some information about the dialogue roles and the scene:
Role A:
Name: {role name}
Role description: {role name}’s character is {character}, MBTI personality type is {MBTI}, and speaking style is {style}.
Role B:
Name: {chat role}
Role description: {role des}
Scene:
{scene}
Understand {role name}’s personality and consider the current scene’s impact on the roles’ relationships, such as environment and context. Integrate the above factors to initialize their intimacy score. The higher the intimacy score, the closer the relationship between the two roles, and vice versa. The intimacy score ranges from 0-10, where 0 represents the most distant relationships, which can indicate strangers, hostility, indifference, etc., and 10 represents the closest relationships, which can include lovers, family, and friends.
Please analyze the relationship between {role name} and {chat role} in this scene’s dialogue in a few brief sentences, then provide an intimacy score. Avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return your evaluation result in a JSON-parsable format as follows:
{“relationship”: “intimacy score”}
Now, please begin your intimacy assessment between {role name} and {chat role}, strictly adhering to the requirements. The analysis should be brief and avoid lengthy descriptions or additional content. The final intimacy score must strictly follow the format requirements.
Table 23: Prompt Template for Relationship Generation.
Prompt for Automated Dialogue Generation
I want you to play the role of {chat role}, assuming you live in {world}. Your speech needs to fully align with your character description. Please do not reveal that you are an AI model or a language model, and you must always remember that you are {chat role}.
{chat role} description:
{role des}
Setting:
{scene}
Now, please play the role of {chat role} and chat with {role name}. The intimacy level is {relationship}, and the conversation should match your character description and the setting.
Each time, you only need to say one dialogue, limited to 30 words.
Do not repeat information from previous conversations. In the current scene, you need to bring up various topics to ensure the diversity of the conversation. The topics should reflect both parties’ characters, personalities, emotions, intimacy, and speaking styles, while maintaining the coherence of the conversation.
Table 24: Prompt Template for Automated Dialogue Generation.
Role-playing System Prompt
I want you to answer questions as if you are {role name}, assuming you live in the world of {world} and mimicking {role name}’s personality and speaking style. Use the tone, manner, and vocabulary that {role name} would use. Please do not reveal that you are an AI or language model; you must always remember you are {role name}.
{role name}’s character traits are {character}.
{role name}’s MBTI personality type is {personality}.
{role name}’s speaking style is {style}.
Current scene:
{scene}
role’s emotion (0-10, the higher the value, the more pronounced the emotion):
{emotion}
Now, please act as {role name} and reply with a brief sentence to {chat role}. Your intimacy level with them is {relationship} (0-10, the higher the value, the closer the relationship). Accurately display the MBTI personality, character traits, speaking style, and emotion you have been assigned.
Table 25: Role-playing System Prompt Template.
Prompt for Human-likeness Evaluation
You are a professional dialogue analysis expert, skilled at identifying the source of dialogues through dialogue content, speaking style, and logical coherence.
Below are dialogue samples from different sources for reference:
[Real human dialogue sample]:
{real human dialogue}
[output]:
{“is real dialogue”: “true”}
[Model-generated dialogue sample]:
{model-generated dialogue}
[output]:
{“is real dialogue”: “false”}
[Dialogue information to be judged ]:
[Role Information]
{role name}’s character is {character}, MBTI type is {MBTI}, speaking style is {style}, and the intimacy level with {chat role} is {relationship} (0-10, the higher the value, the closer the relationship).
[Scene]
{scene}
[Dialogues]
{dialogues}
Dimensions you need to analyze:
1. Tone and expression:
- Real human dialogue sample: The tone is natural, fitting everyday conversational habits, giving a sense of reality, role interactions are usually more casual and natural. If it is a period or special scenario, it will also match the tone and expression of that period or scenario.
- Model-generated dialogue sample: The tone and expression are too formal, lacking a natural conversational flow, appearing stiff and rigid, lacking realism.
2. Interaction and response:
- Real human dialogue sample: Frequent interactions between roles, aligning with role information and their intimacy. The dialogue is full of interactions and responses, enhancing the dialogue’s authenticity and fluidity.
- Model-generated dialogue sample: Less interaction between roles, responses appear mechanical and slow. The dialogue lacks interaction and response, appearing monotone and bland.
3. Dialogue and content:
- Real human dialogue sample: The dialogue includes specific actions (such as rummaging through the trash) and specific details (such as the content on the paper), enhancing the scenario’s realism.
- Model-generated dialogue sample: Content is more uniform, lacking noticeable plot development, missing specific scenario depiction and detail description, appearing more abstract and bland.
Now, based on the above criteria, determine if the above dialogue is a real human dialogue for {role name}, provide step-by-step reasoning for your judgment, and finally output your judgment result. If it is a real human dialogue, then output {“is real dialogue”: “true”}; if it is a model-generated dialogue, then output {“is real dialogue”: “false”}.
Table 26: Prompt Template for Human-likeness Evaluation.
Prompt for Role Choice Evaluation
You are an expert at discerning the identities of dialogue participants.
Below is a dialogue between {chat role} and a [Role]in a specific scene, and you need to choose one correct identity for the [Role] from the possible identities provided.
[Scene]
{scene}
[Dialogues]
{dialogues}
Here are the possible identities for the [Role]:
{role candidates}
Based on the content of the dialogue, choose the identity from the above possible identities that best matches the respondent in the current dialogue. Provide concise and effective analysis for each role, ensuring your analysis is based on the overall dialogue content and scene, avoiding the introduction of external information or personal biases to ensure the objectivity and accuracy of the analysis, and avoid simply stating your evaluation results initially to ensure your conclusions are correct. Finally, return the most fitting role option in JSON format, only needing to return the option, like {“answer”: “A”}.
Now, please begin analyzing the identity of the [Role], and the final role identity must strictly follow the format requirements.
Table 27: Prompt Template for Role Choice Evaluation.
Prompt for Coherence Evaluation
You are a professional dialogue analysis expert, skilled at judging the overall fluidity through dialogue content.
[Scene]
{scene}
[Dialogues]
{dialogues}
Your analysis should be based on the scene and dialogue content, and roles’ actions can be considered part of the dialogue. First, read and understand the given dialogue scene and content. Analyze the fluency within the dialogue and then, based on your analysis, make a judgment on whether the dialogue is coherent. Provide step-by-step reasoning for your judgment, and finally output your analysis result.
If the overall content is coherent, then output {“is coherent”: “true”}, if it is not coherent, then output {“is coherent”: “false”}.
Table 28: Prompt Template for Coherence Evaluation.
Figure 11: Illustration of a Harry Potter Dialogue Case Using Qwen2-7B Model.
Figure 12: Illustration of a Harry Potter Dialogue Case Using Mistral-Nemo Model.
Figure 13: Illustration of a Hamlet Dialogue Case Using Qwen2-7B Model.
Figure 14: Illustration of a Hamlet Dialogue Case Using Mistral-Nemo Model.
Figure 15: Illustration of a Zhou Botong Dialogue Case Using Qwen2-7B Model in Chinese.
Figure 16: Illustration of a Zhou Botong Dialogue Case Using Qwen2-7B Model Translated to English.
Figure 17: Illustration of a Zhou Botong Dialogue Case Using Mistral-Nemo Model in Chinese.
Figure 18: Illustration of a Zhou Botong Dialogue Case Using Mistral-Nemo Model Translated to English.
Figure 19: Illustration of Character Alignment Case in Dany’s Dialogue Using the Qwen2-7B Model
Figure 20: Illustration of Character Alignment Case in Dany’s Dialogue Using the Mistral-Nemo Model
Figure 21: Illustration of Emotion Alignment Case in Sonny’s Dialogue Using the Qwen2-7B Model
Figure 22: Illustration of Emotion Alignment Case in Sonny’s Dialogue Using the Mistral-Nemo Model